OpenStudy (anonymous):

log3 8 = x how do i find x

OpenStudy (anonymous):

is this log base 3 of 8?

OpenStudy (anonymous):

if so you will need a calculator.

OpenStudy (anonymous):

yes i don't know what to type in the calculator for some reason i have forgotten

OpenStudy (anonymous):

you are trying so solve \[3^x=8\] and since 8 is not an integral power of x you need a calculator. it is \[\frac{ln(8)}{ln(3)}\] or \[\frac{log(8)}{log(3)}\]

OpenStudy (anonymous):

thank you so so so so so much i don't know how i forgot that so fast

OpenStudy (anonymous):

i meant since 8 is not an obvious power of 3.

OpenStudy (anonymous):

welcome.

OpenStudy (anonymous):

u got time for one more

OpenStudy (anonymous):

I DIDNT GOT IT ANY ONE CAN HELP ME???????????????

OpenStudy (anonymous):

kya hua mian??